OpenStudy (anonymous):

what are two numbers whose difference is 48 and their quotient is 7? how do you do this

OpenStudy (lgbasallote):

two numbers whose difference is 48 so x - y = 48 quotient is 7 so x/y = 7 does that help?

OpenStudy (anonymous):

so am i trying to find two numbers that subtract and equal 48 and the same two numbers divide and equal 7

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

is that even possible 7 doesn't even go into 48

OpenStudy (lgbasallote):

it is possible..

OpenStudy (lgbasallote):

there may be a chance the numbers are fractions though

OpenStudy (lgbasallote):

you should use algebra to be sure

OpenStudy (anonymous):

ive been trying i cant figure it out it is very frustrating

OpenStudy (lgbasallote):

okay..i gave you x/y = 7 can you rewrite this equation so that it shows x = ?

OpenStudy (anonymous):

what?

OpenStudy (lgbasallote):

do you get how i got x - y = 48 and x/y = 7?

OpenStudy (anonymous):

yes i get that, i don't understand how two numbers subtract = 48 and those same two numbers divided = 7 i have subtracted many numbers that = 48 but when you divide them they don't = 7

OpenStudy (lgbasallote):

stop trying trial and error...this is solveable by algebra

OpenStudy (lgbasallote):

so you have x/y = 7 do you know how to express this in terms of x?

OpenStudy (anonymous):

no

OpenStudy (lgbasallote):

do you know how to cross multiply?

OpenStudy (anonymous):

48 x 7

OpenStudy (lgbasallote):

no... at x/y = 7

OpenStudy (lgbasallote):

\[\huge \frac xy = 7\] cross multiply

OpenStudy (anonymous):

i have no idea

OpenStudy (lgbasallote):

multiply y to 7 what do you get?

OpenStudy (anonymous):

7y

OpenStudy (lgbasallote):

right so x = 7y

OpenStudy (lgbasallote):

now substitute x = 7y into x - y = 48

OpenStudy (anonymous):

x-7y=48

OpenStudy (lgbasallote):

no... 7y - y = 48 x = 7y so you substitute 7y into x does that make snese?

OpenStudy (anonymous):

ok so 6y = 48

OpenStudy (lgbasallote):

right. now solve for y

OpenStudy (anonymous):

y=8?

OpenStudy (lgbasallote):

yes

OpenStudy (lgbasallote):

so now solve for x x - 8 = 48

OpenStudy (anonymous):

x=56

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so there are your numbers.. 56 and 8

OpenStudy (lgbasallote):

anyway...i need to go now so good luck

OpenStudy (anonymous):

i am jumping up and down cheering for myself thank you so much

OpenStudy (lgbasallote):

welcome